OpenStudy (anonymous):

I need help!!!!!!!!! -(1+7x)-6(-7-x)=36 Please Help

OpenStudy (anonymous):

@ladiesman217?

OpenStudy (ladiesman217):

what are you solving for

OpenStudy (anonymous):

For x.

OpenStudy (anonymous):

-1-7x+42+6x=36 41-36=7x-6x 5=x

OpenStudy (ladiesman217):

x=5

OpenStudy (anonymous):

thankyou;)

OpenStudy (ladiesman217):

your welcome